OpenStudy (anonymous):

At the city museum, child admission is$5.50 and adult admission is $9.10. On Wednesday, 134 tickets were sold for a total sales of $1032.20. How many adult tickets were sold that day?

OpenStudy (texaschic101):

a + c = 134----> c = 134 - a 9.10a + 5.50c = 1032.20 now sub 134 - a in for c into the second equation..then just solve for a

OpenStudy (anonymous):

But what's a?

OpenStudy (texaschic101):

a = adult tickets and c = childrens tickets let me set it up for you... 9.10a + 5.50(134 - a) = 1032.20 can you solve it now ?

OpenStudy (anonymous):

a=82?

OpenStudy (anonymous):

Can you help me set up another problem?

OpenStudy (anonymous):

Jim's Coffee Shop makes a blend that is a mixture of two types of coffee. Type A coffee costs Jim $4.25 per pound, and type B coffee costs $5.30 per pound. This month's blend used three times as many pounds of type B coffee as type A, for a total cost of $705.25. How many pounds of type A coffee were used?

OpenStudy (texaschic101):

9.10a + 5.5(134 - a) = 1032.20 9.10a + 737 - 5.5a = 1032.20 9.10a - 5.5a = 1032.20 - 737 3.6a = 295.20 a = 295.20/3.6 a = 82...you got this one right ================= 4.25a + 5.30b = 705.25 3a = b 4.25a + 5.3(3a) = 705.25 4.25a + 15.9a = 705.25 20.15a = 705.25 a = 705.25/20.15 a = 35 3a = b 3(35) = b 105 = b check.. 4.25a + 5.30b = 705.25 4.25(35) + 5.30(105) = 705.25 148.75 + 556.50 = 705.25 705.25 = 705.25 (correct) type A coffee = 35....type B coffee = 105